# neurips_2025.bib

### v1

@@ Abstract
Recent vision-language-action (VLA) models built on pretrained vision-language models (VLMs) have demonstrated strong generalization across manipulation tasks. However, they remain constrained by a single-frame observation paradigm and cannot fully benefit from the motion information offered by aggregated multi-frame historical observations, as the large vision-language backbone introduces substantial computational cost and inference latency. We propose CronusVLA, a unified framework that extends single-frame VLA models to the multi-frame paradigm through an efficient post-training stage. CronusVLA comprises three key components: (1) single-frame pretraining on large-scale embodied datasets with autoregressive action tokens prediction, which establishes an embodied vision-language foundation; (2) multi-frame encoding, adapting the prediction of vision-language backbones from discrete action tokens to motion features during post-training, and aggregating motion features from historical frames into a feature chunking; (3) cross-frame decoding, which maps the feature chunking to accurate actions via a shared decoder with cross-attention. By reducing redundant token computation and caching past motion features, CronusVLA achieves efficient inference. As an application of motion features, we further propose an action adaptation mechanism based on feature-action retrieval to improve model performance during finetuning. CronusVLA achieves state-of-the-art performance on SimplerEnv with 70.9% success rate, and 12.7% improvement over OpenVLA on LIBERO. Real-world Franka experiments also show the strong performance and robustness. Project website.
Figure 1: CronusVLA is a multi-frame modeling framework that begins with single-frame pretraining on large-scale manipulation datasets. During post-training on high-quality cross-embodiment datasets, multi-frame encoding and cross-frame decoding are included. The action adaptation is conducted during finetuning. Evaluations in simulation and the real world demonstrate that CronusVLA substantially outperforms prior baselines, with fast inference and long-horizon compatibility.
@@ 1 Introduction
Current low-level policies [diffusionpolicy, zeng2024learning, tian2024robokeygen, perceiver, kalashnikov2018scalable, 3d_vla, huang2025roboground] trained on expert demonstrations have achieved strong task-specific performance, yet their generalization capabilities remain limited due to constrained model capacity and the absence of broad pretraining. The rise of vision-language models (VLMs) [paligemma2, prismatic, llava, phi3] has paved the way for general vision-language-action (VLA) models by offering powerful backbones and pretrained vision-language representations. Recent VLA methods [openvla, RT2, spatialvla, Tracevla] primarily adapt advanced VLMs on large-scale heterogeneous manipulation datasets [oxe, droid, rh20t] by re-engineering the tokenizer, while others [pi0, robovlms, cogact, tinyvla, openvla-oft] draw inspiration from low-level policy designs, incorporating techniques like specialized heads and action chunks for better performance.
Prior low-level policies [ACT, gr1, gr2, seer] have shown that leveraging multi-frame historical inputs across the temporal dimension considerably improves performance, as motion information from multi-frame observations helps determine the current execution phase and effectively resolve ambiguities. However, most existing VLA models [RT2, palme, openvla, spatialvla, Tracevla], built on the single-frame paradigm of VLMs, are typically trained with only a single current observation and instruction. Directly feeding multiple historical observations, as commonly done in low-level policies, poses two key challenges: (1) the self-attention computation of vision-language backbones (i.e. the language model and vision encoders) cost scales quadratically with the length of input tokens, hindering large-scale embodied pretraining; (2) redundant visual tokens considerably degrade inference speed, limiting the practicality of such models in real-world manipulation. TraceVLA [Tracevla] integrates motion information by adding an additional visual-tracing image to OpenVLA [openvla], which may not fully exploit the pretrained capabilities, as OpenVLA and the original VLMs [prismatic] are both rarely trained with such traces.
@@
To facilitate efficient multi-frame modeling, we propose CronusVLA, a general framework for multi-frame training and inference, which is adapted from a single-frame VLA model, as illustrated in Figure 1. The approach comprises three key components: (1) Single-frame Pretraining: We first train a basic single-frame VLA model using standard autoregressive prediction over discrete action tokens, enabling efficient utilization of large-scale heterogeneous embodied datasets and establishing an embodied vision-language foundation. (2) Multi-frame Encoding: We augment the basic single-frame VLA model with learnable motion features and post-train it on high-quality cross-embodiment datasets. This post-training effectively adapts the prediction of vision-language backbones from token-level outputs to feature-level. Aggregating motion features from multiple historical frames into a feature chunking ensures the transition of the model from single-frame awareness to multi-frame. (3) Cross-frame Decoding: the feature chunking is modulated to balance current and past information and then decoded by a cross-frame decoder. This process forms a unified mapping from multi-frame input to future actions and progressively shifts the core of action prediction from autoregressive token prediction to the integration of multiple motion features. We further employ multi-frame regularization, which enhances training convergence capability by limiting the influence of past frames and ensures that only the current frame updates the vision-language backbone.
The core insight of our approach: per-frame feature primarily captures spatial information from a single observation, while by aggregating motion features across multiple time steps, motion patterns naturally emerge. The cross-frame decoder exploits this aggregation to build a unified representation that encodes both spatial structure and temporal dynamics. Our approach offers two advantages in terms of efficiency: (1) Fast inference: Each motion feature inherits motion information from multiple discrete action tokens, enabling flexible action prediction in once forward process without autoregressive token generation. Additionally, past motion features can be cached to avoid redundant computation in the vision-language backbone. These designs enable substantial inference speedups over prior VLA models, even with the added complexity of multi-frame modeling (see Figure 1). (2) Long-horizon compatibility: The cross-frame decoder employs a cross-attention architecture that enables comprehensive decoding while mitigating the typical computational burden of long sequences, resulting in near-constant inference speed regardless of sequence length. Given the strong correlation between motion feature chunkings and executed actions, we further introduce an action adaptation mechanism based on feature-action retrieval. This mechanism retrieves standard actions as informative priors to guide action prediction, ensuring more accurate action prediction.
@@
  • We propose CronusVLA, a general end-to-end framework that extends VLA models to the multi-frame paradigm. Based on single-frame pretraining, CronusVLA unifies multi-frame encoding and cross-frame decoding for action prediction, enabling scalable manipulation learning.
  • We design a multi-frame post-training strategy that aggregates motion information across frames and decodes it using a cross-frame decoder. This approach enables efficient action prediction while also supporting fast inference and long-horizon compatibility. An action adaptation mechanism is further introduced to provide action prior, resulting in considerable performance gains.
  • We conduct extensive experiments across three embodiments and diverse manipulation tasks in both the simulation and real world. CronusVLA achieves state-of-the-art performance on the simulation benchmark SimplerEnv [simplerenv] with an average 70.9% success rate and achieves a 12.7% overall improvement over OpenVLA on LIBERO [Libero] benchmark. CronusVLA also demonstrates strong performance and robustness across real-world simple and long-horizon tasks with Franka platform.
@@ 2 Related Works
Vision-Language-Action models. Current VLA models usually integrate action generation based on the framework of VLMs [prismatic, llava, phi3, gao2025genmanip]. By employing an action tokenizer, RT-2 [RT2] discretizes 7D actions and employs the PaLI-X [paliX] for autoregressive prediction, while OpenVLA also tokenizes actions and trains the Prismatic [prismatic] on the OXE dataset [oxe]. LLaRVA [llarva] directly outputs 2D waypoints and actions following the LLaVA [llava] framework. SpatialVLA [spatialvla] unifies the action space of various robots via their adaptive action grids. 3D-VLA [3d_vla] and Magma [magma] unify action prediction and multimodal embodied tasks within a single model. These methods aim to minimally adapt VLMs into a general-purpose manipulation policy by treating actions as discrete tokens. Instead, the remaining works [tinyvla, roboflamingo, robovlms, pi0] abandon the discrete formulation and do not aim to preserve the core paradigm of VLMs. They augment the original VLMs with additional action heads [LSTM, diffusion, DiT, flowmatching], meanwhile, train embodiment capabilities and continuous action prediction from scratch. Our method focuses on how to transfer a single-frame, discretely pretrained embodied VLA model to the multi-frame setting with continuous action prediction.
@@
Multi-frame modeling for robotic manipulation. Most early VLAs [openvla, palme, RT2, spatialvla, cogact] treat each action prediction as a temporally independent decision, and are also trained in a single-frame manner. In contrast, low-level policies [RT1, diffusionpolicy, ACT, vip] incorporate multi-frame modeling by processing multiple images simultaneously based on their lightweight architectures. Other policies [gr1, gr2, seer] are pre-trained on large-scale video generation tasks [howto100m, ego4d, droid] by interleaving multi-step action and multi-frame image prediction, and effectively learn multi-frame awareness from diverse training data. However, directly applying this strategy to large-scale VLAs introduces considerable computational overhead. To address these limitations, prior work has made preliminary attempts. TraceVLA [Tracevla] uses visual prompting to draw the multi-frame past trace on the current observation. The most related works to ours are RoboFlamingo [roboflamingo] and RoboVLMs [robovlms], which primarily adopt memory-based heads to model temporal relations across frames. They forcibly extend single-frame pretrained VLMs to multi-frame action prediction by training embodiment capabilities from scratch, akin to standard policy learning. However, they overlook the benefits of effectively transferring a single-frame pretrained VLA model to the multi-frame paradigm, both in terms of training efficiency and performance. In contrast, our method builds on a single-frame pretrained VLA model and explicitly establishes multi-frame capabilities in an additional decoder during post-training, which retains the single-frame perception meanwhile enabling multi-frame modeling.
@@
Figure 2: Overview of CronusVLA framework. (a) illustrates the training of the basic single-frame VLA. By duplicating the model weights, we perform multi-frame post-training as shown in (b), where multi-frame modeling is achieved by aggregating motion features from several preceding frames in a cross-frame decoder. In (c), the upper section details the architecture of the decoder, while the lower section shows how the action adapter generates reference actions to serve as prior information.
@@ 3 Methodology
We present the details of our training strategy and model design in this section. In Section 3.1, we describe the single-frame training process. In Section 3.2, we introduce the multi-frame encoding approach and motion features. Cross-frame decoding are shown in Section 3.3, focusing on the cross-frame decoder and multi-frame regularization. Finally, Section 3.4 details the action adaptation mechanism. The overview is illustrated in Figure 2.
@@ 3.1 Single-frame Pretraining
As illustrated in Figure 2 (a), our first step is to establish a foundational vision-language backbone. Off-the-shelf pretrained VLMs [prismatic, qwen2vl] are adapted into our basic single-frame VLA model by learning diverse manipulation demonstrations $D_{i}=\left(I_{t},a_{t},l\right)_{t=0}^{T_{i}}$ [droid, oxe], where $T_{i}$ is the length of episode $i$. Here, $l$ is the language instruction, $I_{t}$ denotes the observation from a one-view camera at step $t$, and $a_{t}\in\mathbb{R}^{n}$ represents the corresponding actions. Discrete action tokens are derived from continuous robot actions $a_{t}$ via the extended action tokenizer, which maps them into 256 bins, and are trained using the next-token prediction objective with token-level cross-entropy loss, following [RT2, openvla]. Given $I_{t}$ and $l$, the model predicts the next-step action tokens and detokenizes them, $a_{t}=\mathrm{VLA}(I_{t},l)$. We observe that the single-frame pretraining effectively transfers the visual perception capabilities of vision encoders [siglip, dinov2] to embodied scenes, which provides an effective vision-language foundation for multi-frame post-training. Meanwhile, it can better maintain the single-frame visual perception and multi-modal understanding learned during general vision-language pretraining.
3.2 Multi-frame Encoding
@@
From discrete action tokens to motion feature chunking. Vision tokens $\left\{v^{i},i\in\left[0,n_{v}\right]\right\}$ and text tokens $\left\{l^{i},i\in\left[0,n_{l}\right]\right\}$ are causally computed in the vision-language backbone of our basic single-frame VLA, it autoregressively predict discrete action tokens by summarizing information from all previous tokens. As shown in Figure 2 (b), instead of generating discrete action tokens $a_{t}$, we introduce learnable motion features ${f}_{t}\in\mathbb{R}^{d}$ in the backbone’s hidden layers as continuous representations. This feature is designed to integrate the pretrained model’s embodied vision-language summarization capability and is computed as ${f}_{t}=\mathrm{VL}(I_{t},l)$. As our basic VLA models are in the single-frame formulation, we introduce feature chunking ${F}_{t}^{M}=\{{f}_{t-M+1},\dots,{f}_{t-1},{f}_{t}\}$ to effectively represent a multi-frame relationship. It is a chunking of historical motion features and can represent multi-frame observations of $M$ steps at the feature-level. During training, we perform multi-frame prediction over $M$ steps by restructuring inputs at the batch level, enabling the vision-language backbone to independently process $B\times M$ single-frame inputs per iteration, where $B$ denotes the original batch size. This yields the motion feature chunking ${F}_{t}^{M}$ for the cross-frame decoder. During inference, we maintain the feature chunking using a first-in, first-out queue mechanism, which considerably accelerates inference by reusing prior vision-language computations.
3.3 Cross-frame Decoding
@@
Cross-frame decoder. The cross-frame decoder performs action prediction by using the multi-frame motion information embedded within the feature chunking ${F}_{t}^{M}$ to get action chunking $a_{t:t+K-1}=\mathrm{Decoder}({F}_{t}^{M})$, as shown in Figure 2 (b~c). Following [DiT], we construct a Transformer-based decoder composed of self-attention network and MLP layers, and train it using a diffusion loss $\mathcal{L}_{\mathrm{diff}}$. To balance the contributions of the current and past motion features in action prediction, we employ a modulator to dynamically modulate the motion features. Specifically, the current motion feature $f_{t}\in\mathbb{R}^{d}$ is divided to match the number of past motion features $M-1$ through $\mathrm{DIV}$ function, and processed together to produce the modulated feature ${Z}_{f}$:
@@
$Z_{f}=\mathrm{Modulator}(F_{t}^{M})=\mathrm{MLP}\left(\{{f}_{t-M+1},\dots,{f}_{t-1}\},\tilde{f}_{t}\right),\text{where }Z_{f}\in\mathbb{R}^{2\cdot(M-1)\times d^{\prime}},$ (2)
@@
where $\mathrm{DIV}$ consists of a dimensionality-expanding MLP followed by a feature-splitting operation. We further adopt a cross-attention mechanism to separate actions and motion features, which enables effective interaction while avoiding increased computational overhead, ensuring the decoder remains scalable to longer horizons. Specifically, $Z_{f}$ is fed into the cross-attention network and mapped to the keys and values, where noised actions $\hat{a}$ serve as queries. Noised actions are iteratively denoised conditioned on $Z_{f}$ for the final action output.
@@
Post-training with multi-frame regularization. We introduce the multi-frame regularization to decouple the vision-language backbone from multi-frame modeling in the decoder, ensuring its training remains consistent with the single-frame paradigm. Specifically, the past motion features $\{{f}_{t-M+1},\dots,{f}_{t-1}\}$ within the feature chunking ${F}_{t}^{M}$ are treated as auxiliary inputs to the decoder, with their influence limited to the decoding part. Their gradient flow is restricted, preventing any updates to the vision-language backbone, and just serve solely as a regularization term to facilitate training. The overall objective changes, where $\mathrm{sg}$ means the stop-gradient operation:
$\{{f}_{t-M+1},\dots,{f}_{t-1}\}=\left\{\mathrm{sg}\big{(}\mathrm{VL}(I_{t-k},l)\big{)}\,\middle|\,k=1,\dots,M{-}1\right\},$ (3)
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\mathrm{diff}}\left(\hat{a}_{t:t+K-1},a_{t:t+K-1}\mid\{{f}_{t-M+1},\dots,{f}_{t-1}\},\tilde{f}_{t}\right),$ (4)
@@
where $\mathrm{VL}$ denotes processed by the vision-language backbone. This method offers two advantages: (1) Extracting past motion features without gradient computation reduces computational and memory overhead, enabling efficient training. (2) Updating on a single-frame basis preserves the pretrained single-frame perceptual capabilities of the backbone network and promotes faster convergence.
3.4 Action Adaptation
Figure 3: An illustration of the retrieval matrix construction.
During finetuning, we employ an action adaptation mechanism for action prediction as shown in Figure 2 (c). We observe that motion feature chunkings ${F}_{t}^{M}$ effectively capture action patterns from expert demonstrations and exhibit strong correlations with future actions. This enables the use of feature chunkings to retrieve standard actions as coarse priors for guiding action prediction. In Figure 3, each multi-frame clip $I_{t-M:t}$ of all expert demonstrations, with $\mathrm{length}=M$, is processed to extract the motion feature chunkings ${F}_{t}^{M}$ and its associated action sequences. The feature chunking ${F}_{t}^{M}$ is then transformed into modulated features $Z_{f}$ via the modulator. Assuming there are $N$ clips, we simply flatten and normalize ${Z}_{f}\in\mathbb{R}^{C\times d^{\prime}}$ into $\hat{{Z}_{f}}\in\mathbb{R}^{D}$, and then form the retrieval matrix $X\in\mathbb{R}^{N\times D}$, which supports a feature-to-action search. The adapter performs action retrieval by computing the cosine similarity of the key (i.e., the current $\hat{{Z}_{f}}$) to all vectors in the retrieval matrix $X$, that is $s=X\hat{Z}_{f}^{\top}$. The top-k most similar entries $\hat{s}$:
$\hat{s},\mathcal{I}=\operatorname{Topk}(s),\quad w=\operatorname{softmax}(\hat{s})\in\mathbb{R}^{N}$ (5)
where $\mathcal{I}\in\mathbb{R}^{k}$ denotes the indices of the top-k matches, $w$ is the normalized similarity weights. The reference action is aggregated from the retrieved entries as:
$\hat{a}_{t:t+K-1}=\begin{cases}{0},&\text{if }\max(\hat{s})<\tau\\ \sum\limits_{i=1}^{k}w_{i}\cdot a^{\mathcal{I}_{i}}_{t:t+K-1},&\text{otherwise},\end{cases}$ (6)
where $\tau$ is a confidence threshold to reject ambiguous retrievals. If similarity is low, $a^{\mathcal{I}_{i}}_{t:t+K-1}$ will be set to zero vectors to indicate uncertainty. Otherwise, it is weighted by corresponding actions to produce the reference actions $\hat{a}_{t:t+K-1}$. Inspired by fusion mechanisms in video [blattmann2023stable] and action generation [robodual] that integrate both priors and noisy information to predict future states, we propose concatenating $\hat{a}_{t:t+K-1}$ with noised actions and projecting them into the decoder’s shared feature space via a linear layer, serving as a initial state to guide final action generation.

### v2

# CronusVLA: Towards Efficient and Robust Manipulation via Multi-Frame Vision-Language-Action Modeling

@@ Abstract
Recent vision-language-action (VLA) models built on pretrained vision-language models (VLMs) have demonstrated strong performance in robotic manipulation. However, these models remain constrained by the single-frame image paradigm and fail to fully leverage the temporal information offered by multi-frame histories, as directly feeding multiple frames into VLM backbones incurs substantial computational overhead and inference latency. We propose CronusVLA, a unified framework that extends single-frame VLA models to the multi-frame paradigm. CronusVLA follows a two-stage process: (1) Single-frame pretraining on large-scale embodied datasets with autoregressive prediction of action tokens, establishing an effective embodied vision-language foundation; (2) Multi-frame post-training, which adapts the prediction of the vision-language backbone from discrete tokens to learnable features, and aggregates historical information via feature chunking. CronusVLA effectively addresses the existing challenges of multi-frame modeling while enhancing performance and observational robustness. To evaluate the robustness under temporal and spatial disturbances, we introduce SimplerEnv-OR, a novel benchmark featuring 24 types of observational disturbances and 120 severity levels. Experiments across three embodiments in simulated and real-world environments demonstrate that CronusVLA achieves leading performance and superior robustness, with a 70.9% success rate on SimplerEnv, a 26.8% improvement over OpenVLA on LIBERO, and the highest robustness score on SimplerEnv-OR. These results highlight the potential of efficient multi-frame adaptation in VLA models for more powerful and robust real-world deployment. Project Website.
Figure 1: CronusVLA is a multi-frame modeling framework that includes single-frame pretraining on large-scale manipulation datasets and multi-frame post-training on cross-embodiment datasets. CronusVLA shows fast inference, high performance in simulation benchmarks and real-world experiments, and better observational robustness.
@@ 1 Introduction
The rise of vision-language models (VLMs) (Karamcheti et al. 2024; Wang et al. 2024) has paved the way for general vision-language-action (VLA) models by offering powerful backbones and pretrained vision-language representations. Recent VLA methods (Kim et al. 2024; Qu et al. 2025; Zheng et al. 2024) primarily adapt advanced VLMs on large-scale heterogeneous manipulation datasets (O’Neill et al. 2024; Khazatsky et al. 2024) by re-engineering the tokenizer, while others (Black et al. 2024; Li et al. 2024a) draw inspiration from low-level policy designs, incorporating techniques like specialized heads and action chunks for better performance.
Low-level policies (Wu et al. 2023; Cheang et al. 2024; Tian et al. 2024) have shown that leveraging multi-frame historical observations (i.e., RGB images captured from a single viewpoint over time) can considerably enhance performance. In fact, multi-frame information offers two notable advantages: (1) motion cues derived from consecutive observations help determine the current execution phase and effectively resolve state ambiguities. (2) actions can be reliably inferred from prior consistent observations, even when the current input is corrupted, demonstrating strong observational robustness during execution.
However, most existing VLA models (Brohan et al. 2023; Kim et al. 2024; Qu et al. 2025), built on the single-frame paradigm of vision-language models (VLMs), are typically trained using only a single current observation. Directly feeding multiple historical observation images into VLA models introduces two major challenges: (1) the self-attention computation in VLM backbones scales quadratically with the number of input tokens, making large-scale embodied pretraining computationally expensive; and (2) redundant visual tokens considerably degrade inference speed, limiting the feasibility of real-world deployment. RoboVLMs (Li et al. 2024c) attempt to extend single-frame pretrained VLMs to the multi-frame paradigm by adopting memory-based LSTMs (Graves and Graves 2012) and training embodiment capabilities from scratch, following standard policy learning paradigms. However, they overlook the potential benefits of efficiently adapting single-frame pretrained models to the multi-frame paradigm, and lack quantitative evaluation of the performance gains brought by such multi-frame capabilities.
@@
To facilitate efficient multi-frame modeling, we propose CronusVLA, a general framework for multi-frame training and inference, as illustrated in Fig.1. The approach comprises two key components: (1) Single-frame Pretraining: We first train a basic single-frame VLA model using standard autoregressive prediction over discrete action tokens, enabling more convenient utilization of large-scale heterogeneous embodied datasets and establishing an effective vision-language foundation. (2) Multi-frame Post-training: We introduce learnable features in the basic single-frame VLA model and conduct post-training on high-quality cross-embodiment datasets. This multi-frame post-training will replace multiple discrete action tokens with continuous learnable features and adapt the prediction of vision-language backbones from single-frame awareness to multi-frame. Extracting motion cues from multiple historical frames into a feature chunking enables effective temporal information aggregation, thereby improving efficiency. We further introduce the feature modulator and multi-frame regularization, which mitigate temporal imbalance and enhance convergence capability by reconstructing the influence of past frames within the model.
Current benchmarks, such as SimplerEnv (Li et al. 2024b) and LIBERO (Liu et al. 2024), primarily focus on evaluating manipulation models across diverse tasks, objects, and scenes. They largely overlook the impact of observational disturbances on manipulation models, which is a critical issue for future real-world applications. We introduce the SimplerEnv-OR (Observational Robustness) benchmark, which enables quantitative evaluation of model robustness beyond standard training data augmentation, considering both temporal and spatial disturbances. It includes 24 types of observational disturbances with over 120 levels of severity, and evaluates performance across 2,300 trials.
In total, our approach offers three advantages: (1) Fast inference. CronusVLA predicts learnable features in a single forward pass without relying on autoregressive prediction, enabling faster action generation over previous VLA models. Cached historical learnable features further eliminate redundant computation during inference, resulting in substantial speed improvements. (2) High performance. Extensive experiments across three embodiments and diverse manipulation tasks in both the simulation and real world are conducted. CronusVLA achieves state-of-the-art performance on the simulation benchmark SimplerEnv with an average 70.9% success rate, a 26.8% overall improvement over OpenVLA on LIBERO. (3) Strong robustness. CronusVLA can maintain strong robustness against temporal and spatial disturbances. In our real-world experiments, it achieves a 72.6% success rate under various challenges involving interference and occlusion. On our SimplerEnv-OR benchmark, the model consistently surpasses previous approaches in both robustness score and success rate.
@@
  • We propose a general framework, CronusVLA, to extend VLA models to a multi-frame paradigm, which includes single-frame pertaining and multi-frame post-training.
  • We propose SimplerEnv-OR, a novel benchmark that enables quantitative evaluation of model robustness under observational disturbances.
  • Extensive experiments have demonstrated CronusVLA’s leading performance and strong observational robustness across simulation and the real world.
@@ 2 Related Works
Vision-Language-Action models. Current VLA models usually integrate action generation into pretrained VLMs (Karamcheti et al. 2024; Abdin et al. 2024). By employing an action tokenizer, RT-2 (Brohan et al. 2023) and OpenVLA discretize 7D actions and generate them via autoregressive prediction. SpatialVLA (Qu et al. 2025) unifies the action space of various robots via adaptive action grids. 3D-VLA (Zhen et al. 2024) and Magma (Yang et al. 2025a) unify action prediction and multimodal embodied tasks within a single model. These methods minimally adapt VLMs into a general manipulation policy by treating actions as discrete tokens. Instead, the remaining works (Wen et al. 2024; Li et al. 2024c; Yang et al. 2025b) abandon the discrete formulation and do not aim to preserve the original paradigm of VLMs. They augment the original VLMs with additional action heads (Graves and Graves 2012; Peebles and Xie 2023), meanwhile, train embodiment capabilities and continuous action prediction from scratch. Our method focuses on transferring a single-frame, discretely pretrained VLA model to the multi-frame paradigm with continuous action prediction.
@@
Multi-frame modeling for manipulation. Most VLAs (Kim et al. 2024; Driess et al. 2023; Brohan et al. 2023) treat each action prediction as a temporally independent decision, and are also trained in a single-frame manner. While low-level policies (Chi et al. 2023; Zhao et al. 2023) incorporate multi-frame modeling by processing multiple images simultaneously based on their lightweight architectures. Other policies (Wu et al. 2023; Tian et al. 2024) are pre-trained on large-scale video datasets (Miech et al. 2019; Grauman et al. 2022) by interleaving multi-step action and multi-frame image prediction. However, directly applying this strategy to large-scale VLAs introduces considerable computational overhead. To address these limitations, Dita (Hou et al. 2025) directly employs multiple RGB images. TraceVLA (Zheng et al. 2024) uses visual prompting to draw past traces on the current observation. RoboFlamingo (Li et al. 2023) and RoboVLMs (Li et al. 2024c) primarily adopt memory-based LSTMs and training embodiment capabilities from scratch to model temporal relations across frames. In contrast, our method builds on a single-frame pretrained VLA model and explicitly establishes multi-frame capabilities during post-training, which retains the single-frame perception meanwhile enabling multi-frame modeling.
@@
Figure 2: Overview of CronusVLA framework. (a) illustrates the single-frame pretraining of the basic single-frame VLA. By duplicating the model weights, we perform multi-frame post-training as shown in (b), where multi-frame modeling is achieved by aggregating learnable features from several preceding frames in a cross-frame decoder. In (c), a queue mechanism is conducted on feature chunking for fast inference. Details of the cross-frame decoder are illustrated in (d).
@@ 3 Methodology
We present the details of our training strategy and model design in this section. In Sec.3.1, we describe the single-frame training process. In Sec.3.2, we detail the multi-frame post-training approach, including the vision-language backbone, the cross-frame decoder, and multi-frame regularization. Finally, Sec.3.3 show our SimplerEnv-OR benchmark. The overview is illustrated in Fig.2.
@@ 3.1 Single-frame Pretraining
As shown in Fig.2 (a), our first step is to establish a basic single-frame VLA model. Off-the-shelf pretrained VLMs (Karamcheti et al. 2024; Wang et al. 2024) are adapted by learning large-scale manipulation demonstrations $D_{i}=\{I_{t},a_{t},l\})_{t=0}^{T_{i}}$, where $T_{i}$ is the length of episode $i$, $l$ is the language instruction, observation $I_{t}$ denotes a RGB image from a single camera at step $t$, and $a_{t}\in\mathbb{R}^{n}$ represents the corresponding actions. Following (Kim et al. 2024), discrete action tokens are derived from continuous actions $a_{t}$ via the extended action tokenizer, which maps actions into 256 bins, and are trained using the next-token prediction objective. Given $I_{t}$ and $l$, the model predicts the next-step action tokens and detokenizes them, $a_{t}=\mathrm{VLA}(I_{t},l)$. We observe that the single-frame pretraining effectively transfers the visual perception capabilities of vision encoders SigLip (Zhai et al. 2023) and Dinov2 (Oquab et al. 2023) to embodied scenes, which provides an effective vision-language foundation for multi-frame post-training. Meanwhile, single-frame pretraining can better preserve single-frame visual perception of VLMs and incur lower training costs on large-scale data, compared to directly pretraining in a multi-frame manner.
3.2 Multi-frame Post-traing
@@
From discrete action tokens to feature chunking. For our basic single-frame VLA, vision tokens $\left\{v^{i},i\in\left[0,n_{v}\right]\right\}$ and text tokens $\left\{l^{i},i\in\left[0,n_{l}\right]\right\}$ are causally computed in the vision-language backbone, which autoregressively predict discrete action tokens by summarizing information from all previous tokens. As shown in Fig.2 (b), during multi-frame post-training, instead of generating discrete action tokens $a_{t}$, we introduce learnable features ${f}_{t}\in\mathbb{R}^{d}$ in the backbone’s hidden layers as continuous representations. This feature is designed to integrate the pretrained model’s embodied vision-language summarization capability and is computed as ${f}_{t}=\mathrm{VL}(I_{t},l)$. All images are still encoded by our vision-language backbone within a single-frame formulation, ensuring compatibility with standard VLM formulations. We introduce feature chunking ${F}_{t}^{M}=\{{f}_{t-M+1},\dots,{f}_{t-1},{f}_{t}\}=f_{t-M+1:t}$ to establish the association between different frames. It is a chunking of historical learnable features and can represent multi-frame images of $M$ steps at the feature-level. During training, we restructure $M$-step image inputs at the batch level, enabling the vision-language backbone to independently process $B\times M$ single-frame inputs per iteration, where $B$ denotes the original batch size. During inference, as shown in Fig.2 (c), we maintain the feature chunking using a first-in, first-out queue mechanism, which ensures fast inference by reusing prior vision-language computations.
@@
Cross-frame decoder. The cross-frame decoder predicts action chunking by decoding the multi-frame information embedded in the feature chunking ${F}_{t}^{M}$, $a_{t:t+K-1}=\mathrm{Decoder}({F}_{t}^{M})$, as shown in Fig.2 (d). We construct a DiT-based decoder composed of self-attention networks and MLP layers, and train it using a diffusion loss $\mathcal{L}$. To balance the contributions of the current and past learnable features, we employ a feature modulator to dynamically modulate the learnable features. Specifically, the current feature $f_{t}\in\mathbb{R}^{d}$ is divided to match the number of past learnable features $f_{t-M+1:t-1}$ through channels splitting ($\mathrm{DIV}$), and then processed by modulator ($\mathrm{MD}$) to get the modulated feature ${Z}_{f}$:
$Z_{f}=\mathrm{MD}(F_{t}^{M})=\mathrm{MLP}\left({f}_{t-M+1:t-1},\tilde{f}_{t}\right)$ (1)
@@
where $\mathrm{DIV}$ consists of a dimensionality-expanding Linear layer followed by a feature-splitting operation, $Z_{f}\in\mathbb{R}^{2\cdot(M-1)\times d^{\prime}}$, $d^{\prime}$ is the hidden dimension of the decoder. We further adopt a cross-attention mechanism to process noised actions and modulated features, which enables effective interaction. Specifically, $Z_{f}$ is fed into the cross-attention network and mapped to the keys and values, where noised actions $\hat{a}$ serve as queries. Noised actions are iteratively denoised conditioned on $Z_{f}$ for the final action output during inference.
@@
Post-training with multi-frame regularization. We introduce the multi-frame regularization to decouple the vision-language backbone from multi-frame modeling of the decoder, ensuring its training logic remains consistent with the single-frame paradigm. Specifically, the past learnable features ${f}_{t-M+1:t-1}$ of the feature chunking ${F}_{t}^{M}$ are treated as auxiliary inputs to the decoder, with their influence limited within the decoder, where $t$ is the execution step. Their gradient flows do not update the vision-language backbone, and serves solely as a regularization term to facilitate training:
$\hat{f}_{t-M+1:t-1}=\left\{\mathrm{sg}\big(\mathrm{VL}(I_{t-k},l)\big),k=1,\dots,M{-}1\right\},$ (3)
$\mathcal{L}=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I}),i}\left[\left\|\hat{\epsilon}^{i}-\epsilon_{\theta}\left(t,\ \hat{f}_{t-M+1:t-1},\ f_{t}\right)\right\|_{2}\right],$ (4)
@@
where $\mathrm{sg}$ means the stop-gradient operation, $\mathrm{VL}$ is the vision-language backbone, $\mathcal{L}$ is the diffusion loss (multi-step denoising), $\hat{\epsilon}^{i}$ is the predicted $i$-step noise, associated with the noised actions $\hat{a}^{i}_{t:t+K-1}$. This method offers two advantages: (1) Extracting past learnable features without gradient computation reduces computational and memory overhead, enabling efficient training. (2) Updating on a single-frame basis preserves the pretrained single-frame perceptual capabilities and promotes faster convergence.
3.3 SimplerEnv-OR Benchmark
Figure 3: An illustration of the SimplerEnv-OR benchmark.
As shown in Fig.3, we introduce the SimplerEnv-OR benchmark. It is designed for quantitatively evaluating the observational robustness of VLAs by simulating different observational disturbances of the camera, which is beyond the simple visual augmentation strategies used in training.
Evaluation settings. SimplerEnv-OR extends the simulation environment of the WidowX Robot Visual Matching (WR-VM) setting in SimplerEnv (Li et al. 2024b), and evaluates models trained on Bridge-v2 (Walke et al. 2023). We consider two disturbance dimensions: spatial and temporal. The spatial dimension introduces visual disturbances in different positions to simulate real-world camera artifacts, including Global (i.e., blurring, jittering, frame dropping, full occlusion), Local (i.e., overexposing, partial occlusions), and Discrete (i.e., noise, impulse). The temporal dimension assesses model robustness under varying disturbance frequencies, including Constant (1:0), Cyclic (1:1), and Sparse (1:3, 1:5, …). Under all disturbance settings, VLAs must complete the same tasks as WR-VM. Details are provided in Appendix.A.
Robustness score. To quantify robustness, we define a robustness score $R\text{-}Score$ based on relative performance maintenance under various disturbances. Let $SR$ denote the average success rate of the original WR-VM tasks, and $SR^{i}$ denote the success rate under the disturbance setting $i$. The robustness score is computed as: $R\text{-}Score^{i}=100*\frac{SR^{i}}{SR}$. Noting that each setting includes 200 to 400 trials, ensuring a stable and reliable evaluation.